OpenStudy (pounce0129):

Use substitution to solve each system of equations y=3x+8 5x+2y=5

OpenStudy (anonymous):

you plug iin the value of y into y of the second equation. 5x + 2(3x + 8) = 5. then you should get 5x + 6x + 16 = 5, and solve it from there.

OpenStudy (anonymous):

5x+2(3x+8) = 5 5x + 6x + 16 = 5 5x + 6 x - 16 = -16 11x = -16 divide both side by 11 x = -16/11

OpenStudy (anonymous):

no, you forgot to minus 16 from 5. the correct answer should be 11x = -11, with x = -1.

OpenStudy (pounce0129):

thanks :)

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

whoooppsss haha thank you for checking, totally forgot!